An asteroid has acquired a net negative chargeof 126 C from being bombarded by thesolar wind over the years, and is currently in equilibrium wherebyit expels electrons at the same rate as it acquires them. How manymore electrons does it have than protons?

a.
1.26 ×10-17
b.
788
c.
756
d.
7.88 ×1020
Answer #1
Given q = 126C
We know the formula is q = ne
Where n is the number of protons
           e is the charge of the proton = 1.6*10-19C
Now 126 = (n)(1.6*10-19C)
          n =78.75*1019protons
            = 7.88 ×1020